# Weighted Ensemble Models Are Strong Continual Learners

### v1

@@ 4.3 Ablation Study
In CL, a primary objective is to balance knowledge retention from previous tasks with the acquisition of new information from current tasks. This study examines the impact of the $\lambda$ hyperparameter used to balance the weight of the models when averaging.
In addition to our method, we also include the aggregation scheme of Wise-FT [75] for comparison. More specifically, we introduce the following additional baselines: (i) WiSE-FT ($\theta_{0}$) employs an averaging strategy between the model fine-tuned on the current task ($\theta_{t}$) and the initial pre-trained model ($\theta_{0}$), following the formula $\theta_{t}=\lambda\theta_{t}+(1-\lambda)\theta_{0}$. This method aligns with the WiSE-FT [75] and concentrates exclusively on the initial pre-training and the current task without integrating knowledge from preceding tasks. (ii) WiSE-FT ($\theta_{t-1}$) adapts Wise-FT to the CL context by incorporating sequential fine-tuning model with the initial pre-trained model $\theta_{0}$. and (iii) CoMA (Ours).
According to Fig. 4, CoMA demonstrates superior performance compared to both versions of Wise-FT across various $\lambda$ settings. This improvement suggests that our method is more effective at incorporating new knowledge while preserving old information. The best performance is achieved with $\lambda\!=\!0.3$, giving the best trade-off between learning new task $t$ (i.e plasticity), while preserving knowledge from previous tasks (i.e stability). For CoMA, the extreme values for $\lambda$ present scenarios similar to Seq FT, we continue fine-tuning the previous/current model if $\lambda\!=\!0$/$\lambda\!=\!1$.
Furthermore, the performance of WiSE-FT ($\theta_{t-1}$) surpasses that of WiSE-FT ($\theta_{0}$), highlighting the benefit of incorporating sequential learning from previous tasks. As $\lambda$ increases, there is an observed enhancement in the performance of WiSE-FT ($\theta_{t-1}$). However, it remains less effective compared to CoMA, this underperformance is attributed to the impact of averaging with the model $\theta_{0}$, which possesses significantly different parameter values due to fine-tuning. Consequently, this averaging process results in suboptimal parameter values as explained in Sec. 3.
In summary, our method, CoMA, effectively maintains a balance between retaining old task information and adapting to new task data. This is achieved by leveraging weight averaging between successive models. The key benefit here is that the parameter values of models at tasks $t$ and $t-1$ do not diverge significantly, ensuring that the averaged parameters remain effective for both tasks [6, 24].
@@ 5 Conclusion
In this work, we have introduced a novel approach, CoFiMA, designed to address the challenge of catastrophic forgetting in the CL setting. This approach builds its grounding on two pillars. First, it leverages linear connectivity, providing a balanced mechanism for retaining prior knowledge while accommodating new information. Second, Fisher information is incorporated to intelligently weigh the averaging of parameters. This refinement allows for the adjustment of each parameter’s value based on its importance, as determined by its Fisher information, thereby effectively reducing catastrophic forgetting.
Our benchmarking on diverse datasets with various PTMs backbones demonstrates that CoFiMA consistently outperforms state-of-the-art CIL methods. Our findings underscore the efficacy of CoFiMA in mitigating forgetting and highlight its versatility across different PTM backbones and benchmark datasets.
Acknowledgements. This paper has been supported by Hi!PARIS Center on Data Analytics and Artificial Intelligence. This work was granted access to the HPC resources of IDRIS under the allocation AD011013860 made by GENCI.
Supplementary Material
In this supplementary material, we provide more details about the experimental results mentioned in the main paper, as well as additional empirical evaluations and discussions. The supplementary material is organized as follows: In Section A, we provide the detailed algorithm for CoFiMA. Section B provides additional details on weight-averaging methods. In Section C, we report extra experiments about CoFiMA with different PTM backbones. Finally, in Section D, we detail the baselines, and datasets used in our experiments.
Appendix A Detailed Algorithm
In this section, we provide the detailed implementation of CoFiMA algorithm (see Alg. 2). We utilize the same notations as in SLCA [82] to describe CoFiMA training process.
CoFiMA algorithm begins with inputs that include a pre-training dataset $D_{pt}$, training datasets $D_{t}$ for tasks $t=1,...,T$, and a network $M_{\theta}(\cdot)$ with parameters $\theta=\{\theta_{rps},\theta_{cls}\}$. The parameters $\theta_{rps}$ are initialized by pre-training on $D_{pt}$, and $\theta_{cls}$ are randomly initialized. For each task, the network is trained using cross-entropy loss on $D_{t}$, applying different learning rates $\alpha$ and $\beta$ to $\theta_{rps}$ and $\theta_{cls}$, respectively, until convergence.
@@ Appendix A Detailed Algorithm
Input: Pre-training dataset $D_{pt}$; training dataset $D_{t}$ for task $t=1,...,T$; model $M_{\theta}(\cdot)=h_{\theta_{cls}}(f_{\theta_{rps}}(\cdot))$ with parameters $\theta=\{\theta_{rps},\theta_{cls}\}$; learning rates $\alpha$ and $\beta$ for $\theta_{rps}$ and $\theta_{cls}$ respectively.
Initialization: Pre-train $\theta_{rps}$ on $D_{pt}$; Initialize $\theta_{cls}$ randomly.
@@
2:       Train network $M_{\theta}$ on $D_{t}$ until convergence:
@@
4:             Update $M_{\theta}$ using cross-entropy loss.
@@
6:       Weight update process:
7:       Compute Fisher Information $F_{t}$ for $\theta_{t}$.
8:       Update weights $\theta^{*}_{t}$ as per Eq.7.
9:       Calculate Fisher Information $F_{t}^{*}$ for $\theta^{*}_{t}$.
10:       Feature collection:
11:       for each class $c\in C_{t}$ do
12:             Collect feature set $F_{c}=[r_{c,1},...,r_{c,N_{c}}]$.
13:             Compute mean $\mu_{c}$ and covariance $\Sigma_{c}$ of $F_{c}$.
14:       end for
@@
20:Train classifier $h_{\theta_{cls}}$:
@@
Implementation Specifics. A pre-trained ViT-B/16 backbone is utilized for our method as in [82]. For approaches that maintain fixed backbones and employ prompting, an Adam optimizer is used [72, 73] as in the original papers; conversely, for baselines that involve updating the entire model, an SGD optimizer is employed, both using a uniform batch size of 128. Our approach employs a learning rate of 0.0001 for the PTM and 0.01 for the classification layer. Empirically, Zhang et al. [82] found that supervised pre-training tends to reach convergence more rapidly than self-supervised pre-training. Thus, for supervised pre-training, we train all baselines for 20 epochs on Split CIFAR-100 and 50 epochs on other datasets. For self-supervised pre-training, a training duration of 90 epochs is applied across all benchmarks.
Appendix B More Details about Baselines
@@
  • Weight-Ensemble [74]: The key idea of this work is that by averaging the weights of multiple models that have been fine-tuned with different hyperparameter configurations, it is possible to improve both the accuracy and robustness of the resulting model. This method is referred to as ”model soups”. One of the main benefits is that, unlike conventional ensemble methods, averaging many models in this way does not incur additional inference or memory costs.
@@
  • WiSE-FT [75] introduces a method for enhancing robustness by ensembling the weights of the zero-shot and fine-tuned models. This approach aims to balance the need for high accuracy on target distributions while preserving the model’s ability to perform well on varied and unforeseen data distributions as follows: $\theta_{\text{WiSE-FT}}=\alpha\theta_{\text{zero-shot}}+(1-\alpha)\theta_{\text{fine-tuned}}$.
@@
  • Exponential Moving Average (EMA) [62]: EMA in deep learning models is a technique used to smooth out data by giving more weight to recent observations. It is often used in training neural networks to stabilize and improve the learning process by averaging the parameters over time: $\theta_{m}=\beta\theta_{m}+(1-\beta\theta_{m-1})$. In our setting, we found that the best performance is achieved by using $\beta=0.999$ on the evaluated datasets.
@@
    CoFiMA differentiates itself from papers using EMA by leveraging Fisher information to adjust the averaging process, prioritizing parameters based on their relevance to preserve knowledge from previous tasks in CL. Furthermore, CoFiMA’averaging process is done after learning completely the task, not after some epochs. In contrast, EMA applies a decay factor to average model weights over time, primarily to stabilize training and improve convergence, without a task-specific focus on mitigating forgetting.
Appendix C Ablation Backbones

### v3

@@ 4.3 Ablation Studies
Balancing the Information of an Old and a New Task. In CL, a primary objective is to balance knowledge retention from previous tasks with the acquisition of new information from current tasks. This study examines the impact of $\lambda$ used to balance the stability/plasticity of the model during averaging.
Figure 3: CoFiMA with various PTMs on CIFAR-100. CoFiMA enhances the results of SLCA.
@@
In addition to our method, we also include an adaptation of the aggregation scheme of WiSE-FT [80] for CL, which we refer to as WiSE-FT-CL. After learning task $t$, stability is achieved via linear interpolation between the pre-trained and the fine-tuned model: $\boldsymbol{\theta}_{t}^{*}=\lambda\boldsymbol{\theta}_{t}+(1-\lambda)\boldsymbol{\theta}_{0}$. For every new task $t$, fine-tuning starts from $\boldsymbol{\theta}_{t-1}^{*}$. This method aligns with WiSE-FT [80] and relies only on the initial pre-training to achieve stability.
According to Fig. 4, CoFiMA demonstrates superior performance compared to WiSE-FT-CL across various $\lambda$ settings. This improvement suggests that our method is more effective at incorporating new knowledge while preserving old information. The best performance is achieved with $\lambda\!=\!0.3$, giving the best trade-off between learning new task $t$ (i.e., plasticity) and preserving knowledge from previous tasks (i.e., stability).
The underperformance of WiSE-FT-CL is attributed to the impact of averaging with the model $\boldsymbol{\theta}_{0}$, which possesses significantly different parameter values than $\boldsymbol{\theta}_{t}$ due to fine-tuning. Consequently, this averaging process results in suboptimal parameter values as explained in Sec. 3. For CoFiMA, higher values for $\lambda$ put more emphasis on the current model at task $t$, thus leading to forgetting previous tasks.
In summary, our method, CoFiMA, effectively maintains a balance between retaining old task information and adapting to new task data. This is achieved by leveraging weight averaging between successive models. The key benefit here is that the parameter values of models at tasks $t$ and $t-1$ do not diverge significantly, ensuring that the averaged parameters remain effective for both tasks [6, 21] (more details in supplementary material).
@@ 5 Conclusion
In this work, we have introduced CoFiMA, the first method based on weight averaging techniques to address catastrophic forgetting in the CIL setting. This approach builds its grounding on two pillars. First, it leverages model averaging, providing a balanced mechanism for retaining prior knowledge while accommodating new information. Second, Fisher information is incorporated to intelligently weigh the averaging of parameters. This refinement allows for the adjustment of each parameter’s value based on its importance, as determined by its Fisher information, thereby effectively reducing catastrophic forgetting.
Our benchmarking on diverse datasets with various PTM backbones demonstrates that CoFiMA consistently outperforms state-of-the-art CIL methods. Our findings underscore the efficacy of CoFiMA in mitigating forgetting and highlight its versatility across different PTM backbones and benchmark datasets.
Acknowledgements
We thank Yasser Benigmim, Khalid Oublal and Ekaterina Iakovleva for helpful discussions and draft feedback. This paper has been supported by the French National Research Agency (ANR) in the framework of its JCJC, and was funded by the European Union’s Horizon Europe research and innovation program under grant agreement No. 101120237 (ELIAS). This research was additionally funded by the EU project AI4TRUST (No.101070190). Furthermore, this research was partially funded by Hi!PARIS Center on Data Analytics and Artificial Intelligence. This work was granted access to the HPC resources of IDRIS under the allocation AD011013860 made by GENCI.
Appendix 0.A More details CoMA/CoFiMA
0.A.1 Detailed Derivations
In the main paper, in Section 3.2, we explain that using $\lambda=\frac{1}{t}$ in Eq. (3) is equivalent to the arithmetic average with uniform weights as in Eq. (2). We substantiate this claim through an inductive proof, detailed in the ensuing paragraphs.
Let us denote by $\bar{\theta}_{t}$ the arithmetic mean of $t$ distinct model parameters $\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots,\boldsymbol{\theta}_{t}$. Commencing with the base case where $t=1$, the equivalence is intuitively apparent as $\bar{\boldsymbol{\theta}}_{1}=\boldsymbol{\theta}_{1}$.
To validate the inductive step, we proceed with the subsequent mathematical derivations:
$\bar{\boldsymbol{\theta}}_{t}$ $=\frac{1}{t}\sum_{k=1}^{t}\boldsymbol{\theta}_{k}$
$=\frac{1}{t}(\boldsymbol{\theta}_{t}+\sum_{k=1}^{t-1}\boldsymbol{\theta}_{k})$
$=\frac{1}{t}\boldsymbol{\theta}_{t}+\frac{t-1}{t}\frac{1}{t-1}\sum_{k=1}^{t-1}\boldsymbol{\theta}_{k}$
$=\frac{1}{t}\boldsymbol{\theta}_{t}+\frac{t-1}{t}\bar{\boldsymbol{\theta}}_{t-1}$
$=\lambda\boldsymbol{\theta}_{t}+(1-\lambda)\bar{\boldsymbol{\theta}}_{t-1},\text{with~{}}\lambda=\frac{1}{t}.$ (8)
0.A.2 Detailed Algorithm
In this section, we provide the detailed implementation of CoFiMA algorithm (see Alg. 1). We utilize the same notations as in SLCA [85] to describe CoFiMA’s training process.
CoFiMA begins with inputs that include a pre-training dataset $D_{pt}$, training datasets $D_{t}$ for tasks $t=1,...,T$, and a network $M_{\theta}(\cdot)$ with parameters ${}{\theta=\{\theta_{rps},\theta_{cls}\}}$. The parameters $\theta_{rps}$ are initialized by pre-training on $D_{pt}$, and $\theta_{cls}$ are randomly initialized. For each task, the network is trained using cross-entropy loss on $D_{t}$, applying different learning rates $\alpha$ and $\beta$ to $\theta_{rps}$ and $\theta_{cls}$, respectively, until convergence.
@@ 0.A.2 Detailed Algorithm
Input: Training dataset $\mathcal{D}_{t}$ for task $t=1,...,T$; network $M_{\varphi}(\cdot)$ with pre-trained parameters $\boldsymbol{\theta}_{0}^{*}$;
@@
2:     ##Train network $M_{\phi}$ on $\mathcal{D}_{t}$ until convergence:
@@
4:         Update $M_{\phi}$ using cross-entropy loss.
@@
6:     ##Feature collection:
7:     for each class $c\in C_{t}$ do
8:         Collect feature set $F_{c}=[r_{c,1},...,r_{c,N_{c}}]$.
9:         Compute mean $\mu_{c}$ and covariance $\Sigma_{c}$ of $F_{c}$.
10:     end for
11:     ##Weight update process:
12:     Compute Fisher Information $\mbox{\bf F}_{t}$ for $\boldsymbol{\theta}_{t}$.
13:     Update weights $\boldsymbol{\theta}^{*}_{t}$ as per Eq.7.
14:     Calculate Fisher Information $\mbox{\bf F}_{t}^{*}$ for $\boldsymbol{\theta}_{t}^{*}$.
@@
20:##Train classifier $h_{\theta_{cls}}$:
@@
Appendix 0.B Comparison with alternative methods
0.B.1 Comparison to EWC (Elastic Weight Consolidation)
Elastic Weight Consolidation (EWC) [26]. EWC leverages the Fisher information matrix as a regularization mechanism within the loss function, penalizing deviations from previously learned tasks’ parameters. The loss function in EWC is defined as:
$\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}_{B}(\boldsymbol{\theta})+\sum_{i}\frac{\lambda}{2}F_{i}(\boldsymbol{\theta}_{i}-\boldsymbol{\theta}^{*}_{A,i})^{2}$ (9)
where $\mathcal{L}_{B}(\boldsymbol{\theta})$ is the loss for the current task B, $\lambda$ represents the trade-off between learning new tasks and retaining old knowledge, $F_{i}$ denotes the Fisher information matrix elements for the previous tasks (that we name A), and $\boldsymbol{\theta}^{*}_{A,i}$ are the optimal parameters for A.
Comparison and Discussion.
  • Utilization of Fisher Information. EWC employs the Fisher information to inform the penalty within the loss function, thereby retaining critical features from previous tasks. Conversely, CoFiMA uses the Fisher information to modulate the parameter averaging, facilitating a balanced integration of knowledge from consecutive tasks.
  • Objective Function Implications. The loss function in EWC is augmented with a regularization term, making the optimization process more complex. In contrast, CoFiMA’s weighted averaging does not alter the loss function, potentially leading to a simpler optimization landscape.
  • Memory and Computational Requirements. EWC necessitates storage of the entire history of Fisher information matrices and parameters, scaling with the number of tasks. CoFiMA, on the other hand, might only require the Fisher information matrix for the current and immediately preceding task, offering a more memory-efficient alternative.
Both EWC and CoFiMA offer innovative solutions to catastrophic forgetting. EWC emphasizes the preservation of performance on past tasks through penalization of parameter shifts, while CoFiMA promotes a dynamic balance across tasks via weighted averaging. The selection between these two approaches depends on the specific requirements of the application and the computational resources at hand.
0.B.2 Comparison to Weight-Averaging Methods
@@
  • Weight-Ensemble [79]: The key idea of this work is that by averaging the weights of multiple models that have been fine-tuned with different hyperparameter configurations, it is possible to improve both the accuracy and robustness of the resulting model. This method is referred to as "model soups". One of the main benefits is that, unlike conventional ensemble methods, averaging many models in this way does not incur additional inference or memory costs.
@@
  • WiSE-FT [80] introduces a method for enhancing robustness by ensembling the weights of the zero-shot and fine-tuned models. This approach aims to balance the need for high accuracy on target distributions while preserving the model’s ability to perform well on varied and unforeseen data distributions as follows: $\boldsymbol{\theta}_{\text{WiSE-FT}}=\alpha\boldsymbol{\theta}_{\text{zero-shot}}+(1-\alpha)\boldsymbol{\theta}_{\text{fine-tuned}}$.
@@
  • Exponential Moving Average (EMA) [66]: EMA in deep learning models is a technique used to smooth out data by giving more weight to recent observations. It is often used in training neural networks to stabilize and improve the learning process by averaging the parameters over time: $\boldsymbol{\theta}_{m}=\beta\boldsymbol{\theta}_{m}+(1-\beta)\boldsymbol{\theta}_{m-1}$. In our setting, we found that the best performance is achieved by using $\beta=0.999$ on the evaluated datasets.
@@
    CoFiMA differentiates itself from works [60, 49, 86] using EMA by leveraging Fisher information to adjust the averaging process, prioritizing parameters based on their relevance to preserve knowledge from previous tasks in CL. Furthermore, CoFiMA’s averaging process is done after learning completely the task, not after some epochs. In contrast, EMA applies a decay factor to average model weights over time, primarily to stabilize training and improve convergence, without a task-specific focus on mitigating forgetting.
Appendix 0.C More Experimental details
We outline the implementation details, additional experimental results, and description of the methods against which our approach is evaluated, and the datasets used in evaluation:
0.C.1 Implementation Specifics
A pre-trained ViT-B/16 backbone is utilized for our method as in [85]. For approaches that maintain fixed backbones and employ prompting, an Adam optimizer is used [77, 78] as in the original papers; conversely, for baselines that involve updating the entire model, an SGD optimizer is employed, both using a uniform batch size of 128. Our approach employs a learning rate of 0.0001 for the PTM and 0.01 for the classification layer. Empirically, Zhang et al. [85] found that supervised pre-training tends to reach convergence more rapidly than self-supervised pre-training. Thus, for supervised pre-training, we train all baselines for 20 epochs on CIFAR-100 and 50 epochs on other datasets. For self-supervised pre-training, a training duration of 90 epochs is applied across all benchmarks.
0.C.2 Analysing Catastrophic Forgetting via Weight Distance
To further evaluate the effectiveness of CoFiMA in continual learning, as discussed in Sec. 3, we examine how CoFiMA achieves a balance between plasticity and stability from a perspective of Weight Distance.
Weight Distance. In addressing the issue of catastrophic forgetting, we prioritize the stability of model parameters as a key indicator of memory retention. This is grounded in the insight that lesser deviation from a model’s previous weight configuration, upon incorporating new data, signals reduced forgetting [67, 75]. In prior studies [54], the weight distance metric is utilized to quantify forgetting. This metric is defined as the $L_{2}$ norm of the difference between the model parameters after learning two consecutive tasks, expressed as:
$\mathcal{W}(\boldsymbol{\theta}_{t},\boldsymbol{\theta}_{t-1})=\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t-1}\|_{2},$ (10)
where $\boldsymbol{\theta}_{t}$ and $\boldsymbol{\theta}_{t-1}$ represent the model parameters at the completion of tasks $t$ and $t-1$, respectively.
A strong advantage of CoMa is that we can demonstrate that, when learning a task $t$, starting from $\boldsymbol{\theta}^{*}_{t-1}$, the update rule of CoMa guarantees a reduced weight distance compared to directly using the parameters $\boldsymbol{\theta}_{t}$ learned on the task $t$:
$\mathcal{W}(\boldsymbol{\theta}^{*}_{t},\boldsymbol{\theta}^{*}_{t-1})$ $=||\boldsymbol{\theta}^{*}_{t}-\boldsymbol{\theta}^{*}_{t-1}||_{2}$
$=||\lambda\boldsymbol{\theta}_{t}+(1-\lambda)\boldsymbol{\theta}^{*}_{t-1}-\boldsymbol{\theta}^{*}_{t-1}||_{2}$
$=||\lambda(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}_{t-1})||_{2}$
$=\lambda\mathcal{W}(\boldsymbol{\theta}_{t},\boldsymbol{\theta}^{*}_{t-1})$
$\leq\mathcal{W}(\boldsymbol{\theta}_{t},\boldsymbol{\theta}^{*}_{t-1})~{}\text{since}~{}\lambda\in[0,1].$ (11)
(a) CIFAR-100
(b) Imagenet-R
Figure 5: Similarity scores using $L_{2}$ norm for CIFAR-100 (a) and Imagenet-R (b) datasets using CoFiMA and SLCA methods.
Note that similar derivations also hold for CoFiMA. To numerically assess the impact of this reduced weight distance, we report the weight distance $\mathcal{W}$ measured after every task in Fig. 5 on two benchmarks. CoFiMA consistently exhibits lower weight distances across all checkpoints when compared to SLCA, indicating a more stable parameter retention throughout the learning process. Specifically, CoFiMA maintains a remarkably low Weight Distance, with values such as 2.00 and 1.67 for CIFAR-100 and CUB-200 respectively, compared to SLCA’s higher values of 3.17 and 2.86 for the same datasets, suggesting a greater degree of parameter shift and, consequently, a higher susceptibility to forgetting.
0.C.3 Ablation on the Effect of $\lambda$
As detailed in the main paper, the CoMA update rule is given by:
$\boldsymbol{\theta}^{*}_{t}=\lambda\boldsymbol{\theta}_{t}+(1-\lambda)\boldsymbol{\theta}^{*}_{t-1},$ (12)
where $\boldsymbol{\theta}^{*}_{t}$ is the parameter after CoMA at task $t$, $\boldsymbol{\theta}_{t}$ is the parameter after training on task $t$, and $\lambda$ is the hyper-parameter that controls the blending of current and past parameters. By considering $\boldsymbol{\theta}^{*}_{1}=\boldsymbol{\theta}_{1}$ for task $t=1$, and iteratively applying Eq (12), we obtain:
$\boldsymbol{\theta}^{*}_{t}=\lambda\boldsymbol{\theta}_{t}+\lambda(1-\lambda)\boldsymbol{\theta}_{t-1}+\lambda(1-\lambda)^{2}\boldsymbol{\theta}_{t-2}+\ldots+\lambda(1-\lambda)^{t-2}\boldsymbol{\theta}_{2}+(1-\lambda)^{t-1}\boldsymbol{\theta}_{1},$ (13)
Figure 6: Assessing performance of CoFiMA using ViT-B/16 (Left) and MoCo-V3 (Right) with different values for $\lambda$. The marker $+$ represents the best performance.
which can be written as follows:
$\boldsymbol{\theta}^{*}_{t}=\lambda\sum_{k=1}^{t}(1-\lambda)^{t-k}\boldsymbol{\theta}_{k}+(1-\lambda)^{t-1}\boldsymbol{\theta}_{1}.$ (14)
The update process for CoMA over multiple tasks unfolds in a recursive fashion, similarly to EMA, but with the distinction that the averaging is done over tasks (or classes in class-incremental learning) rather than iterations within a task. Eq. (14) shows that each past task’s parameters are weighted by an exponentially decreasing factor (for a positive $\lambda<1$) and the distance (in terms of tasks) from the current task $t$. This ensures that more recent tasks have a greater influence on the current parameter set than older tasks, which aligns with the goals of class-incremental learning where the model should adapt to new classes while still remembering the old ones.
Fig. 6 presents the ablation study results for the CoFiMA method, focusing on the impact of the hyperparameter $\lambda$ on model performance across CIFAR-100, CUB-200, ImageNet-R, and Cars-196 datasets. The parameter $\lambda$ regulates the balance between stability and plasticity—retaining previous knowledge versus adapting to new tasks. For ViT-B/16, performance on CIFAR-100 and CUB-200 improves with increasing $\lambda$ to a point, then levels off or drops slightly, indicating an optimal $\lambda$ range for balancing new and old learnings. However, ImageNet-R and Cars-196 show a decline in performance beyond a certain $\lambda$ value, suggesting a need for greater focus on preserving past knowledge. Conversely, with MoCo-V3, CIFAR-100 benefits significantly as $\lambda$ approaches 1, highlighting the advantage of prioritizing recent learning. CUB-200 shows a positive trend up to a certain $\lambda$, after which performance drops, while Cars-196 remains stable across $\lambda$ values, indicating a different balance requirement. ImageNet-R performance worsens with higher $\lambda$, underscoring the importance of maintaining past knowledge for this dataset.
0.C.4 Ablation Backbones